# RE-RAG: Improving Open-Domain QA Performance and Interpretability with Relevance Estimator in Retrieval-Augmented Generation

### v1

@@ Abstract
Retrieval-augmented generation (RAG) framework is showing state-of-the-art performance on open-domain question answering tasks by referencing external knowledge. However, the RAG system faces challenges with performance degradation when it is fed contexts of low relevance or when the relative relevance among the input contexts is inaccurately assessed. In this work, we propose a RE-RAG framework that injects an explicit context relevance estimator (RE) into the RAG system. RE-RAG re-evaluates the retrieved contexts with the proposed context RE and passes the more relevant contexts along with their measure importance to the generator. To train context RE, we propose an unsupervised learning method, which does not utilize any labeled document ranking data to train the context RE. To examine the efficacy of RE-RAG, we examine its performance on Natural Questions and TriviaQA datasets. RE-RAG achieves on-par performance compared to the FiD variants while utilizing fewer contexts (0.25x). We show that the proposed context RE, which was trained with the T5 model, is also applicable to RAG with LLMs(ChatGPT) by improving the performance on NQ (+6.4EM) and TQA (+2.8EM), respecitvely. Lastly, we display that RE can add interpretability to RAG framework as RE score highly correlates with the RE-RAG accuracy. Consequently, RE can be utilized to filter out unanswerable scenarios where context does not contain answers with 38.9%-51.3% accuracy just by examining a set of retrieved contexts.
@@ 1 Introduction
In recent years, the retrieval augmented generation (RAG) framework has shown promising progress in natural language generation, specifically on knowledge-intensive tasks Lewis et al. (2020b). This approach enhances the model’s faithfulness and reliability by leveraging not only limited parametric memory but also additional non-parametric knowledge Lewis et al. (2020b); Luo et al. (2023). In particular, Open-domain question answering (ODQA) is a knowledge-intensive question-answering task where the task requires a model to provide an answer based on factual information when no specified context is provided along with the question. The RAG framework has shown great success in the ODQA problem and motivated many new research endeavors under the RAG framework.
One of the most prominent models among the RAG variants is the Fusion-in-Decoder (FiD) (Izacard and Grave, 2021b) which showed generally higher performance than original RAG. In turn, since the proposal of FiD, most of the research endeavors focused on the FiD architecture (Izacard and Grave, 2021a; Jiang et al., 2022; Fajcik et al., 2021; Asai et al., 2022). However, despite its superior performance, FiD is hard to interpret as it relies on hundreds of documents through soft cross-attention. Futhermore, FiD is incompatible with black-box LLMs such as GPT and PaLM (Brown et al. (2020), Chowdhery et al. (2023)) that do not disclose their parameters.
On the other hand, RAG, although generally lower-performing than FiD, is more interpretable because it generates answers by each context and marginalizes them. Additionally, RAG decoding steps, which marginalizes answers obtained by providing the generator with integrated questions and contexts, can be easily integrated with recent LLMs (Yang et al. (2023), Xu et al. (2023)). For these reasons, this work revisits the classic RAG framework to leverage its advantages, interpretability and applicability to LLMs, with a modification that can uplift the performance at the level of FiD.
In this work, we propose the RE-RAG framework, which extends the RAG with a context relevance estimator (RE) that re-ranks the retrieved context and provides a precise relevance measure. Our RE-RAG framework retains the interpretable decoding structure of the existing RAG, while achieving higher performance through a context relevance estimator.
The main contributions of our work are as follows:
@@
We propose a new framework RE-RAG by expanding RAG with a Relevance Estimator (RE). We further suggest a training method that can train RE without any labeled data on question-context compatibility.
@@
We demonstrate that RE-RAG, enhanced with RE, significantly improves upon the existing RAG and achieves performance on par with FiD that utilizes many more (4x) contexts.
@@
We show that RE can improve LLMs such as GPT, even when it was trained on a much smaller language model.
4.
We explore methods to filter out low-relevance context in advance by having the RE pre-evaluate the context set before passing it to the generator.
@@ 2 Method
In this section, after reviewing the basic RAG framework, we present the RE-RAG model combined with our relevance estimator.
2.1 Basic RAG overview
@@ 2.2 RE-RAG framework
The retriever similarity score is trained to achieve high recall when retrieving multiple contexts, however, it was not initially designed to provide fine-grained relevancy score for aiding RAG generation steps in eq.(2). To address this issue, we propose a context relevance estimator (RE) that re-ranks contexts and provides precise relevance scores to the generator.
@@
Context relevance estimator (RE) measures the relevance between a question and context. We utilize a similar architecture to Nogueira et al. (2020) which utilizes a sequence-to-sequence model as a passage reranker.
Our context RE receives the same input of question and context as the generator, but is trained to generate a classification token ("true" or "false") based on the relevance of the context to the input question. We normalize the probability of generating "true" and "false" tokens to get the final probability of generating the classification token. The obtained probability of a "true" token can independently be an indicator of the relevance of a single context to a given question. When comparing between multiple contexts, the "true" token probability can be converted to logit and used as the relevance score of the retrieved context.
@@ 2.3 Joint training of RE-RAG
Second, to obtain a learning signal for training the context relevance estimator, we calculate the log-likelihood loss of the generator per retrieved context and compute its distribution across contexts as follows:
@@
We calculate the KL-divergence loss between the probability distributions of the generator and the context RE, and use this loss to train the model.
Lastly, in addition to applying a training loss on the probability of generating the classification token, we need to set an additional loss to prevent the context RE from generating tokens other than the classification token. To do this, we utilize the additional loss as the sum of the probability of context RE of generating all tokens other than classification token.
@@ 3 Experimental Setup
| 223M | 223M | 49.9 | 68.2 | 25 | |
| 770M | 223M | 51.4 | 69.5 | 25 | |
| 770M | 770M | 54.0 | 70.2 | 25 |
@@ 3.1 Dataset
We evaluate our performance on two open-domain QA datasets:Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017). To train and evaluate our model, we utilize the context datasets retrieved for each question from NQ and TQA, as used in FiD-KD (Izacard and Grave, 2021a) and Akari (Asai et al., 2022). The dataset includes the top 20 training contexts, while the dev and test sets contain the top 100 contexts retrieved by the retriever. We used 20 contexts for training and the top-25 contexts extracted by the context RE from the top-100 retrieved contexts for inference.
@@ 3.3 Baseline
We investigate whether the performance of RE-RAG is competitive with that of the FiD (Izacard and Grave, 2021b)-based system. FiD has achieved excellent performance on the Question-Answering task, and the FiD-based application system also outperforms the RAG (Lewis et al., 2020b)-based system on the QA task. Most of the models under comparison involve additional training of the retriever (Izacard and Grave, 2021a) or improvements to the retrieve system (Jiang et al., 2022). Therefore, we use the FiD-KD improved retriever’s dataset for baseline comparisons with other models.
@@ 3.4 Model
The two components of our framework, context RE and the generator, utilize the T5 model (Raffel et al., 2020). We utilize the T5-base, T5-large models, and explore three different model sizes depending on the combination of the two models.
@@ 4 Experiment Results
We investigate the QA performance of the RAG system with our newly proposed context relevance estimator (RE). In addition to the QA performance of the whole system, we also examine the performance of the context RE independently.
@@ 4.1 Main Results
The overall accuracy of our system on the two datasets we evaluated, NQ and TQA, is shown in Table 1. Compared to basic RAG, our system RE-RAG shows better performance despite having the same number of total parameters. Our proposed context relevance estimator (RE) leverages the RAG system while providing a more accurate measure of the relevance between question and context, improving the overall reliability of the system. Despite using top-20 contexts as training dataset and top-25 contexts for inference, our model shows competitive performance compared to FiD-KD (Izacard and Grave, 2021a)and ReAtt (Jiang et al., 2022), which used top-100 contexts for training and inference. The context RE enhances the system’s performance, independent of the generator model’s parameter size.
@@
4.2 Effect of using less document
@@
Table 2 shows the performance of our proposed RE-RAG’s context RE as a reranker. Table 2 presents the performance of re-ranking using retriever’s similarity score and context RE’s relevance score in the top-100 context of retriever. For the Recall@k metric, we use the retrieval accuracy used by DPR (Karpukhin et al., 2020), FiD-KD (Izacard and Grave, 2021a), and ColbertQA (Khattab et al., 2021). Although the comparison retriever has been enhanced through knowledge distillation methods using FiD attention scores, our proposed context RE still demonstrated superior performance. In particular, the performance improvement of the context RE over the retriever becomes more pronounced as the number of contexts decreases. This suggests that the proposed context RE is more effective when the number of contexts that can be fed into the generator is limited.
We examined how QA performance changes when inferring answers using fewer documents. Figure 2 shows that our proposed system performs more robustly when the number of available contexts decreases compared to FiD-KD. In particular, the performance degradation is limited for both models up to 10 contexts, but the difference increases when the number of utilized contexts decreases dramatically.

### v2

@@ Abstract
The Retrieval Augmented Generation (RAG) framework utilizes a combination of parametric knowledge and external knowledge to demonstrate state-of-the-art performance on open-domain question answering tasks. However, the RAG framework suffers from performance degradation when the query is accompanied by irrelevant contexts. In this work, we propose the RE-RAG framework, which introduces a relevance estimator (RE) that not only provides relative relevance between contexts as previous rerankers did, but also provides confidence, which can be used to classify whether given context is useful for answering the given question. We propose a weakly supervised method for training the RE simply utilizing question-answer data without any labels for correct contexts. We show that RE trained with a small generator (sLM) can not only improve the sLM fine-tuned together with RE but also improve previously unreferenced large language models (LLMs). Furthermore, we investigate new decoding strategies that utilize the proposed confidence measured by RE such as choosing to let the user know that it is “unanswerable” to answer the question given the retrieved contexts or choosing to rely on LLM’s parametric knowledge rather than unrelated contexts.
@@ 1 Introduction
In recent years, the retrieval augmented generation framework has shown promising progress in natural language generation, specifically on knowledge-intensive tasks. This approach has been studied in many forms, from traditional RAG Lewis et al. (2020b), which aggregates answers from multiple contexts using document relevance scores as weights, to approaches like RALM Ram et al. (2023), which simply utilizes concatenated context as an in-context learning approach for large-language models (LLMs). Retrieval augmented generation enhances the model’s faithfulness and reliability by leveraging nonparametric knowledge on top of parametric knowledge Luo et al. (2023). In particular, the RAG framework has the advantage of being easily adaptable to modern LLMs Brown et al. (2020); Touvron et al. (2023). These advantages have sparked a significant amount of new research Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) focused on the RAG framework.
Despite the great potential of the retrieval augmented generation framework, if the language model is provided with contexts that are not relevant to the query, it will be distracted by these inappropriate contexts, negatively affecting the accuracy of the answers Yoran et al. (2023). While retrievers or re-rankers in existing research have been effective at measuring the relative ranking across contexts to a query, these modules often fail to determine whether top-ranked contexts are actually relevant to the query or not. Furthermore, if a precise relevance score is not used in the traditional RAG framework, it can cause problems such as directing attention to documents that are less likely to answer the query.
In this work, we propose the RE-RAG framework, which extends traditional RAG by incorporating a relevance estimator (RE) to simultaneously measure the precise relative relevance between retrieved contexts and evaluate their confidence, which can be used to classify whether given context is useful for answering the given question. By more accurately measuring the relative relevance between contexts, RE computes precise relevance scores suitable for weighted aggregated answers in the traditional RAG framework and also acts as an efficient reranker. RE trained on a small generator (sLM) not only benefits sLM fine-tuned together with RE but can also be separated and applied to LLMs as well, benefiting both.
By explicitly classifying whether the context is useful for answering the query, the confidence of context measured by RE provides various decoding strategies. If the retrieved context set is irrelevant, we can choose to classify the query as “unanswerable”, while maintaining most of the accuracy for the answerable set. Additionally, if a low-confidence context set is retrieved, which will likely result in wrong answers by parroting the context as is Jia and Liang (2017), we can instead selectively leverage the LLM’s parametric knowledge to improve answer accuracy in most cases.
The main contributions of our work are:
@@
We propose a new framework called RE-RAG by adding an external Relevance Estimator (RE) module. We further suggest a weak supervision training method that can train RE without explicit labeled data on question-context compatibility. (§2.2)
@@
We demonstrate that RE-RAG, enhanced with RE, significantly improves upon the existing RAG. Addtionally, we show that RE trained on a small language model can improve the answer performance of LLMs. (§4.1)
@@
We propose to use the confidence level of the context set measured by RE to answer “unanswerable” for unanswerable context sets with minimal negative effects, or to complement LLM’s parametric knowledge. (§5.1)
@@ 2 Method
In this section, after reviewing the traditional RAG framework, we present the RE-RAG model combined with our relevance estimator.
2.1 Traditional RAG overview
@@ 2.2 RE-RAG framework
The retriever similarity score is trained to achieve high recall when retrieving multiple contexts, however, it was not initially designed to provide fine-grained relevancy score for aiding RAG generation steps in eq.(2). To address this issue, we propose a relevance estimator (RE) that re-ranks contexts and provides precise relevance scores to the generator.
@@
Relevance estimator (RE) measures the relevance between a question and context. We utilize a similar architecture to Nogueira et al. (2020) which utilizes a sequence-to-sequence model as a passage reranker.
Our RE receives the same input of question and context as the generator, but is trained to generate a classification token ("true" or "false") based on the relevance of the context to the input question. We normalize the probability of generating "true" and "false" tokens to get the final probability of generating the classification token. The obtained probability of a "true" token can independently be an indicator of the relevance of a single context to a given question. When comparing between multiple contexts, the "true" token probability can be converted to logit and used as the relevance score of the retrieved context.
@@ 2.3 Joint training of RE-RAG
Second, to obtain a learning signal for training the relevance estimator, we calculate the log-likelihood loss of the generator per retrieved context and compute its distribution across contexts as follows:
@@
We calculate the KL-divergence loss between the probability distributions of the generator and the RE, and use this loss to train the model.
Lastly, in addition to applying a training loss on the probability of generating the classification token, we need to set an additional loss to prevent the RE from generating tokens other than the classification token. To do this, we utilize the additional loss as the sum of the probability of RE of generating all tokens other than classification token.
@@
| 220M | 220M | 49.9 | 68.2 | 25 | |
| 770M | 220M | 51.4 | 69.5 | 25 | |
| 770M | 770M | 54.0 | 70.2 | 25 | |
| Large language models () | |||||
| (Asai et al., 2023) | - | 7B | - | 66.4 | 5 |
| (Asai et al., 2023) | - | 13B | - | 69.3 | 5 |
| 770M | 7B | 45.7 | 67.1 | 5 | |
| 770M | 13B | 46.6 | 70.8 | 5 | |
| RA-DIT (Lin et al., 2023) | - | 65B | 43.9 | 75.1 | 10 |
| - | 70B | 38.1(40.7) | 63.5(66.3) | 10 | |
| - | 70B | 39.5(46.8) | 68.1(72.1) | 10 | |
| - | 175B | 42.9(45.9) | 69.0(70.7) | 10 | |
| (Shi et al., 2023) | - | 175B | 45.5 | 77.3 | 10 |
| 770M | 8B | 49.6 | 73.0 | 10 | |
| 770M | 70B | 48.0 | 72.4 | 10 | |
| 770M | 70B | 50.8 | 75.5 | 10 | |
| 770M | 175B | 49.3 | 72.6 | 10 | |
3 Experimental Setup
@@ 3.1 Dataset
We evaluate our performance on two open-domain QA datasets:Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017). To train and evaluate our model, we utilize the context datasets retrieved for each question from NQ and TQA, as used in FiD-KD (Izacard and Grave, 2021a) and Akari (Asai et al., 2022). The dataset includes the top-20 training contexts, while the dev and test sets contain the top-100 contexts retrieved by the retriever. We used 20 contexts for training and the top-25 contexts extracted by the RE from the top-100 retrieved contexts for inference.
@@ 3.3 Baseline
We investigate whether the performance of RE-RAG is competitive with that of the FiD (Izacard and Grave, 2021b)-based system. FiD has achieved excellent performance on the Question-Answering task, and the FiD-based application system also outperforms the RAG (Lewis et al., 2020b)-based system on the QA task.
We consider an additional baseline to compare the performance of RE when applied to LLMs. We compare the performance of RE and FiD-KD retriever when applied to LLMs. When applying the FiD-KD retriever to LLMs, we compared two methods: traditional RAG, which uses the retriever similarity score to perform answer marginalization, and RALM, which concatenates all context. Furthermore, we compare our performance with other studies Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) that have implemented RAG in LLMs.
@@ 3.4 Model
The two components of our framework, RE and the generator, utilize the T5 model (Raffel et al., 2020). We utilize the T5-base, T5-large models, and explore three different model sizes depending on the combination of the two models.
Additionally, we utilize Llama2 (7B, 13B, 70B), Llama3 111https://github.com/meta-llama/llama3 (8B, 70B), and ChatGPT (“gpt-3.5-turbo-0125” version) as generators to assess if RE brings performance improvements when applied to LLMs. In our experiments, the LLMs used as generators are not fine-tuned for the downstream task.
@@ 4 Experiment Results
We investigate the QA performance of the RAG system with our newly proposed relevance estimator (RE). In addition to the QA performance of the whole system, we also examine the performance of the RE independently.
@@ 4.1 Main Results
The overall accuracy of our system on the two datasets (NQ and TQA) is shown in Table 1. Compared to the traditional RAG, our system, RE-RAG, performs better despite having the same total number of parameters. Our proposed RE improves the reliability of the RAG system by more accurately measuring the relevance between question and context. Our model performed competitively with models based on FiD structures(Izacard and Grave, 2021a; Jiang et al., 2022; Fajcik et al., 2021).
The accuracy of the RE module when applied to Large Language Models (LLMs) is shown at the bottom of Table 1. We only included the RAG-based model in our comparison because the FiD-based model is not applicable to LLMs due to structural differences. The RE module outperforms the FiD-KD retriever when applied to LLMs. When the RE module is applied to Llama2, it surpasses the Self-RAG, where the LMs themselves inspect the retrieved context and generated answers. In TQA, REPLUG with Codex scores slightly higher. The performance of TQA seems to depend more on the generator model than NQ (see Figure 2 for a related discussion), and we believe that this is the reason for the performance difference with Codex. Our model performs better on NQ, which is a more knowledge intensive task.
@@
| Dataset | Model | Recall | Precision | F1 |
| NQ | FiD-KD | 73.2 | 21.9 | 33.7 |
| 10.3 | 31.0 | 15.5 | ||
| 51.3 | 33.9 | 40.9 | ||
| 45.9 | 38.3 | 41.7 | ||
| TQA | FiD-KD | 64.3 | 24.5 | 35.5 |
| 27.2 | 34.2 | 30.3 | ||
| 38.9 | 46.7 | 42.5 | ||
| 39.0 | 43.2 | 41.0 |
4.2 Performance of RE as a reranker and unanswerable set classifier
@@
Table 2 shows the performance of our proposed RE-RAG’s RE as a reranker. For the Recall@k metric, we use the retrieval accuracy used by DPR (Karpukhin et al., 2020), FiD-KD (Izacard and Grave, 2021a), and ColbertQA (Khattab et al., 2021). Although the comparison retriever has been enhanced through knowledge distillation methods using FiD attention scores, our proposed RE still demonstrated superior performance. In particular, RE performs better as the number of contexts decreases, which means that RE is more efficient when there are fewer contexts to utilize.
Table 3 shows the performance of the context relevance estimator (RE) as a “unanswerable” set classifier. “unanswerable” set means that the context set of the top-25 contexts does not contain a gold answer in any context. For classification, we used the cosine similarity score of the hidden representation of the question and context for retriever and the probability of generating a “true” token by the model for RE and MonoT5 (Nogueira et al., 2020). For the optimal threshold, we searched for the value that maximizes F1 score in steps of 0.1 from 0.5 to 0.9 at dev set.
Our RE showed better “unanswerable” set classification performance than FiD-KD retriever or MonoT5 based on F1 score. Looking at the detailed performance, we found that the retriever performed better for recall, but the RE performed better for precision. This is because the retriever classified a large number of context sets as all “unanswerable” sets, while our proposed RE showed a good balance between precision and recall.
| Dataset | Model | Score | Answerable context set | |
| O | X | |||
| NQ | FiD-KD | 58.3 32.7 | 73.4 | |
| RE | 58.3 54.9 | 51.3 | ||
| FiD-KD | 61.5 34.9 | 71.3 | ||
| RE | 61.5 57.9 | 45.9 | ||
| TQA | FiD-KD | 78.7 51.2 | 63.5 | |
| RE | 78.7 77.0 | 38.9 | ||
| FiD-KD | 80.4 51.6 | 62.7 | ||
| RE | 80.4 77.9 | 39.0 | ||
5 Analysis
5.1 Exploring decoding strategies in low confidence context sets
In this section, we review two strategies that can be used when a context set with a low confidence score is retrieved. The confidence score for a context set is determined using the maximum value of the “true” token probability computed by RE for the contexts within the set. We examine the strategy of answering “unanswerable” when a low confidence context set is returned in a small Language Model (sLM), where parametric knowledge is scarce. Additionally, we examine the strategy of directly utilizing parametric knowledge in Large Language Models (LLMs), where parametric knowledge is abundant.
Classify as “unanswerable” Table 4 shows the change in accuracy after letting the model respond with “unanswerable” when the retrieved context set has low confidence. For the confidence threshold value that determines whether the model should respond with “unanswerable”, we chose the value that optimizes the classification performance as determined in Table 3. We evaluate the accuracy by dividing the entire test set into answerable sets, which contain at least one gold answer in the context set, and unanswerable sets, which contain none.
Our RE model shows relatively minor accuracy loss on the answerable set when responding with “unanswerable” for context sets measured with low confidence, but gains significant ability on the unanswerable set. In contrast, the FiD-KD retriever loses a substantial amount of accuracy on the answerable set when it responds with “unanswerable” for low-confidence context sets, resulting in a larger negative effect compared to our model. If we want to preserve the answerable set accuracy of the FiD-KD retriever, its ability to classify “unanswerable” is significantly reduced compared to RE (see Appendix E).
| P-Generator | R-Generator | NQ | TQA |
| (NQ: 31.1/TQA: 64.3) | 46.2 45.9(-0.3) | 68.0 69.3(+1.3) | |
| 47.3 46.5(-0.8) | 71.5 72.1(+0.6) | ||
| 48.0 46.9(-1.1) | 72.4 72.9(+0.5) | ||
| (NQ: 41.3/TQA: 75.1) | 49.6 49.8(+0.2) | 73.0 75.4(+2.4) | |
| 50.8 50.8(-) | 75.5 76.7(+1.2) | ||
| ChatGPT (NQ: 37.7/TQA: 72.0) | ChatGPT | 49.3 49.3(-) | 72.6 73.6(+1.0) |
Selectively using parametric knowledge We further explore how we can effectively utilize the rich parametric knowledge of LLMs. When the confidence of the retrieved context is low, we examine a mixed strategy that optionally bypasses the context and relies solely on the parametric knowledge of the largest model to generate the correct answer. For the confidence threshold value that determines whether the model should answer using only parametric knowledge, we selected the value that optimizes classification performance as determined in Table 3. For each type of model, we utilize the one with the largest number of parameters as the parametric knowledge base.
Table 5 shows the change in accuracy when decoding the answer using the mixed strategy. In most cases, our strategy achieves accuracy gains in TQA without significant losses in NQ, except in cases where parametric knowledge is particularly scarce, such as in NQ on Llama2. NQ is a more knowledge-intensive task compared to TQA, where there is less benefit from utilizing parametric knowledge.
When parametric knowledge can be used effectively, the mixed strategy achieves larger gains in smaller models, and the performance gap narrows compared to larger models. Figure 2 illustrates the relationship between confidence score and accuracy by model size. At high confidence scores on the TQA dataset, small size models achieve similar accuracy to large size models. At low confidence scores, the difference in performance between small and large models becomes more pronounced. When using small size models, higher efficiency can be achieved by utilizing retrieval augmented generation only when a high confidence context set is retrieved, and selectively leveraging the parametric knowledge of large size models when a low confidence context set is retrieved.
| Model | NQ | TQA |
| Baseline | 39.5 | 54.9 |
| Baseline w/ RE score | 43.1 | 60.1 |
| Baseline w/ RE rerank | 46.8 | 63.9 |
| Baseline w/ RE rerank, score | 49.6 | 67.8 |
| 49.9 | 68.2 |
5.2 Ablation Study
Effectiveness of RE We perform an ablation study to investigate the effectiveness of the added RE in RE-RAG. The effect of our proposed RE is twofold. First, it performs better re-ranking than the retriever, selecting more accurate context and passing it to the generator. Second, it calculates a more accurate relevance score than retriever’s similarity score and uses it in the answer marginalization process. In Table 6, the performance of methods with each component of the RE added is presented, using a model that was trained with only the T5-base generator, after removing the RE, as the baseline.
We construct the following experiment to isolate the two effects. First, we apply the top 25 contexts from retriever and their similarity scores to the baseline model. Next, there are the top-25 contexts from the retriever with the RE’s score applied (RE score) and the top-25 contexts from the RE with the retriever’s similarity score applied (RE rerank). Finally, we compare the performance of applying the RE’s top-25 contexts and score to the baseline model (RE rerank, score).
Both effects of the RE are found to be significant in improving the performance of the baseline model. This shows that not only the quality of the context input to the generator plays an important role, but also the score, which means the importance of each context.
| Model | NQ | TQA |
| Baseline | 0.435 | 0.561 |
| - normalization | 0.0005 | 0.0002 |
Remove training components We investigate the impact of removing the regularization process in eq.(3) on the classification performance of RE while training on the model. Table 7 shows how the “true” token probability level output by the RE changes when the normalization process is removed. It can be seen that when the normalization process is removed, RE can only perform the function of re-ranking but loses the function of measuring confidence. This is because the normalization process allows the model to adjust its output strictly between “true” and “false” tokens.
Table 8 shows the difference in EM scores on the dev set when is removed from the training process. We observed that removing from the training process decreases answer performance. We believe that contributes to achieving more optimal performance by using loss information from generator to directly propagate the relative importance of contexts to the RE.
6 Related Works
Previous research has shown that the performance of Question Answering systems can be improved by utilizing external knowledge about questions (Chen et al., 2017). Methods for more accurate retrieval of external knowledge Karpukhin et al. (2020); Khattab et al. (2021); Gao and Callan (2022) have been studied to make these systems more efficient. In open-domain QA, models that extract and use answers from retrieved documents have been studied Karpukhin et al. (2020); Khattab et al. (2021); Cheng et al. (2021), but studies that utilize generative models such as T5 (Raffel et al., 2020) or BART (Lewis et al., 2020a) have become more common Lewis et al. (2020b); Izacard and Grave (2021b). RAG and FiD achieved powerful performance in open-domain QA using different methods. Subsequently, models Izacard and Grave (2021a); Fajcik et al. (2021); Singh et al. (2021); Jiang et al. (2022) that leverage and improve upon the structural advantages of FiD have been proposed. For Atlas (Izacard et al., 2022), state-of-the-art performance was achieved through an improved retriever (Izacard et al., 2021) and scaling up the model. In the case of RAG, there is a study that improved performance by introducing a BERT (Devlin et al., 2019)-based reranker (Glass et al., 2022), but it utilized additional data and high-quality label data when training the reranker.
Recently, large language models (LLMs) such as GPT (Brown et al., 2020) and Llama (Touvron et al., 2023), which have been developed in recent years, face limitations with FiD methods that require encoded data. Consequently, research on RAG models, which can directly input context, has received renewed attention. Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) These approaches have achieved performance improvements by training a retriever, which can also be applied to LLMs, or by performing the review of questions and context within the model itself.
| Model | NQ | TQA |
| Baseline | 49.1 | 67.8 |
| - | 48.0 | 66.7 |
7 Conclusion
We propose the RE-RAG framework, which extends traditional RAG by incorporating RE that can measure the relative relevance and confidence of contexts. We demonstrate that the RE-RAG framework can enhance the performance of traditional RAG. We show that the RE module, as a detachable component, can be combined with modern large language models (LLMs) to improve their performance. Furthermore, we exploree some decoding strategies that leverage the confidence information measured by the RE module to either answer “unanswerable” or selectively utilize the parametric knowledge of the LLMs when a low confidence context set is retrieved. We hope that our research will inspire the exploration of various additional modules for retrieval-augmented generation.
8 Limitation
Our research has primarily focused on improving answer performance in single-hop QA tasks. We have not sufficiently verified the effectiveness of our proposed framework in multi-hop QA tasks. We believe that in the future, we can explore whether the RE-RAG framework can be extended to multi-hop QA.
In our work, we explored a decoding strategy that measures with confidence whether a context is truly useful for a query and classifies low confidence contexts as unanswerable. However, a truly unanswerable query is one where the query cannot be adequately answered even when utilizing the model’s parametric knowledge. We believe that future research needs to be conducted to detect whether the parametric knowledge has knowledge that can adequately answer the query in order to finally classify the unanswerable problem.
Appendix A Dataset Statistics
| Dataset | Train | Dev | Test |
| Natural Questions | 79,168 | 8,757 | 3,610 |
| TriviaQA | 78,785 | 8,837 | 11,313 |